Write the null and alternative hypothesis of given problem: ( 10 points)


    1. It is claimed that the average weight of a bag of biscuit is 250 grams with a standard deviation of 20.5 grams. Would you agree to this claim if a random sample of 50 bags of biscuits showed an average weight of 240 grams, using a 5% level of significance.



1
Expert's answer
2021-09-28T23:52:46-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=250"

"H_1:\mu\not=250"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{240-250}{20.5/\sqrt{50}}\approx-3.4493"

Since it is observed that "|z| = 3.449 > 1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-3.4493)\approx0.000562," and since "p=0.0006<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "250," at the "\alpha = 0.05" significance level.